# The emergent dynamics of double-folded randomly branching ring polymers

### v1

@@ Abstract
The statistics of randomly branching double-folded ring polymers are relevant to the secondary structure of RNA, the large-scale branching of plectonemic DNA (and thus bacterial chromosomes), the conformations of single-ring polymers migrating through an array of obstacles, as well as to the conformational statistics of eukaryotic chromosomes and melts of crumpled, non-concatenated ring polymers. Double-folded rings fall into different dynamical universality classes depending on whether the random tree-like graphs underlying the double-folding are quenched or annealed, and whether the trees can undergo unhindered Brownian motion in their spatial embedding. Locally, one can distinguish (i) repton-like mass transport around a fixed tree, (ii) the spontaneous creation and deletion of side branches, and (iii) displacements of tree node, where complementary ring segments diffuse together in space. Here we employ dynamic Monte Carlo simulations of a suitable elastic lattice polymer model of double-folded, randomly branching ring polymers to explore the mesoscopic dynamics that emerge from different combinations of the above local modes in three different systems: ideal non-interacting rings, self-avoiding rings, and rings in the melt state. We observe the expected scaling regimes for ring reptation, the dynamics of double-folded rings in an array of obstacles, and Rouse-like tree dynamics as limiting cases. Of particular interest, the monomer mean-square displacements of $g_{1}\sim t^{0.4}$ observed for crumpled rings with $\nu=1/3$ are similar to the subdiffusive regime observed in bacterial chromosomes. In our analysis, we focus on the question to which extent contributions of different local dynamical modes to the emergent dynamics are simply additive. Notably, we reveal a non-trivial acceleration of the dynamics of interacting rings, when all three types of local motion are present. In the melt case, the asymptotic ring center-of-mass diffusion is dominated by the contribution arising from coupling of the ring-in-an-array-of-obstacles dynamics with the tree dynamics. This contribution scales inversely with the ring weight and is compatible with a scenario in which constraint release restores a Rouse-like dynamics.
pacs:
Valid PACS appear here
preprint: APS/123-QED
I introduction
@@
Large ring polymers exhibit surprisingly rich behavior. When diffusing through an array of obstacles, they naturally adopt double-folded conformations, forming randomly branching polymer structures or trees [1, 2, 3, 4, 5, 6] (Fig. 1). Interestingly, double-folded tree-like ring conformations have broad significance in biological systems. Examples include plectonemes in supercoiled DNA molecules [7, 8, 9] and, consequently, their role in bacterial chromosome structuring [10, 11, 8, 12], large RNA molecules [13, 14, 15], and the arrangement of DNA in eukaryotic cells during interphase, which bears striking similarities to the behavior of melts of non-concatenated ring polymers [16, 17, 18].
@@
Figure 1: (a) An example of a (tightly wrapped) double-folded ring polymer and (b) the corresponding tree on a trigonal lattice. Small loops represent reptons (bonds of zero length, see Section III), for which adjacent monomers along the ring occupy identical lattice sites.
@@
Double-folded rings fall into different dynamical universality classes depending on whether the random tree-like graphs underlying the double-folding are quenched or annealed [19], and whether the trees can undergo unhindered Brownian motion in their spatial embedding. Specifically, one can identify three distinct types of local dynamics:
@@
    Reptons transporting ring mass along a fixed tree [20, 9].
@@
    Creation or retraction of side-branches where either (i) one of the two double-folded ring segments double-folds onto itself and branches away from the current tree or where (ii) the ring segments wrapping a side branch are retracted into the rest of the tree in the inverse process.
@@
    Brownian motion of tree nodes maintaining the tree connectivity and the secondary structure of the double-folded rings [21].
The effect of repeated local modifications (or “moves” in the Monte Carlo (MC) language we will adopt throughout the article) is not simply additive to the resulting dynamics of the ring and tree conformations. Instead, an emergent dynamics typically arises because the ring “monomers” and tree “nodes” are not independent degrees of freedom but are coupled through (i) the ring connectivity, (ii) the constraint of double-folding restricting ring conformations to those with underlying tree structures, (iii) additional Hamiltonian terms governing interactions between monomers or nodes, branching activity, or chain stiffness, and (iv) local topological constraints arising from the mutual impenetrability of ring and tree backbones.
This coupling is not a particularity of rings. In linear chains, the connectivity between successive monomers gives rise to the characteristic Rouse intrachain dynamics, while topological constraints are at the origin of the reptation dynamics in the tube model [22]. Dynamical analogues can be formulated for rings. In particular, a well-understood limit is the “ring-in-an-array-of-obstacles” (RiAO) dynamics [2, 6, 23], which arises from the coupling of local mass transport along the tree and the creation and deletion of side branches. However, understanding the general case remains a challenge, more particularly in the melt case, where the “obstacles” represent co-localized ring segments that are not fixed in space but instead have dynamics of their own [23]. The purpose of the present paper is to pave the way for solving this challenge by introducing:
@@
  • a procedure to analyze the independency of contributions from different dynamic modes to the experimentally observable monomer and center-of-mass mean-square displacements.
The manuscript is structured as follows: In Section II we introduce the relevant observables and associated exponents, along with a brief summary of the theoretical background. In Section III, we introduce our model and the corresponding dynamic MC scheme to investigate its dynamical properties. In particular, this scheme extends our previous elastic lattice model for simulating tightly double-folded ring polymers, where only single monomers could move [24], by incorporating an additional tree node move akin to diffusive dynamics. We then explain how to systematically and in a fully controlled manner combine the different types of monomer and tree node moves. Finally, we introduce the systems studied: ideal non-interacting double-folded rings, isolated self-avoiding double-folded rings, and melts of double-folded rings. In the Results Section IV, we briefly examine the well-understood static ring properties of these systems, which also serves to validate that the ensembles generated by different combinations of local dynamics remain identical as long as the system conditions are unchanged. We then focus on comparing the diffusion properties of both monomers and the center of mass across different dynamical schemes and available theoretical results. Notably, in melts, we find that the mixture of tree node and ring monomer moves leads to a non-trivial acceleration of the emergent dynamics. The subsequent Discussion Section V explores the functional form and amplitude of the dynamics arising from the coupling in a context that deals more generally with the analysis of data from mixed dynamical schemes and the question of how to identify and distinguish non-trivial coupling effects from situations where the contributions of different local modes to the emergent dynamics are independent. Finally, we provide a brief conclusion in Section.VI.
II Theoretical Background
II.1 Static Properties
The behavior of a double-folded ring polymer [1, 2, 3, 25, 6, 5, 26] can be analyzed through three structural levels [21] (Fig. 1). The primary structure is defined through the connectivity of the monomers along the ring. The secondary structure corresponds to the double folding. It can be represented by mapping the ring onto a graph with the connectivity of the tree. The tertiary structure is defined by the positions of the ring monomers and tree nodes within the embedding (three-dimensional) space. A small set of exponents describes how the expectation values of typical observables scale with the ring weight $N$ (the number of monomers):
@@ II.1 Static Properties
where $L$ is the mean tree contour distance (the shortest distance on the tree) between tree nodes, $R_{g}$ is the radius of gyration of the ring [24]. For ideal, non-interacting systems, the exponents are $\rho=1/2$ and $\nu=1/4$ [27, 28]. For interacting systems, the only known exact result is the value $\nu=1/2$ for self-avoiding trees in $d=3$ [29]. Otherwise, Flory theory [30, 31, 19, 4, 32] predicts $\rho=2/3$ for self-avoiding trees, and $\nu=1/3$ and $\rho=5/9$ for trees in the melt state.
II.2 Dynamical Properties
@@
On time scales where the rings have moved beyond their own size, their CM exhibit diffusive motion: $g_{1}(t)=g_{3}(t)\sim t$, and $g_{2}(t)=2\langle R^{2}_{g}\rangle$. On shorter time scales, the internal dynamics can display a range of intermediate scaling regimes, whose understanding is important to comprehend the diffusion constant in the asymptotic regime. A summary of the scaling exponents for $g_{1}(t)$, $g_{3}(t)$, and the diffusion constant $D$ in the known limiting cases discussed below are given in Table 1.
II.2.1 Slithering dynamics of tightly wrapped rings around fixed trees
For local mass transport via reptons along a fixed tree [33, 22, 20, 9, 34], the monomer motion in the embedding space is given by:
@@ II.2.1 Slithering dynamics of tightly wrapped rings around fixed trees
where $\langle\delta s^{2}(t)\rangle$ denotes the one-dimensional monomer MSD along the tree. Several regimes must then be distinguished. First, for the shortest time scales, $\langle\delta s^{2}(t)\rangle\sim t$, so that:
@@
At longer time scales, the ring connectivity gives rise to a one-dimensional Rouse dynamics typical of a linear polymer, $\langle\delta s^{2}(t)\rangle\sim t^{1/2}$, so that:
@@
Beyond the Rouse time, $\tau_{R}\sim N^{2}$ where $g_{1}(\tau_{R})\sim N^{\nu}\sim\sqrt{\langle R_{g}^{2}\rangle}$, the dynamics crosses over to the reptation regime where all monomers diffuse coherently around the tree. Again $\langle\delta s^{2}(t)\rangle\sim t$, so that:
@@
II.2.2 Rings in an array of fixed obstacles
@@
The combination of repton moves with the local creation or retraction of side branches gives rise to the characteristic motion of rings in an array of fixed obstacles [2, 6] suppressing the sideways motion of tree nodes. Specifically, $g_{1}(t)\sim t^{\frac{2\nu}{2+\rho}}$ and $g_{3}(t)\sim t^{\frac{2\nu+1}{2+\rho}}$ in the intermediate regime. Moreover, the CM diffusion coefficient is found to scale as $D\sim N^{-2-\rho+2\nu}$ [6].
@@
Table 1: Summary of the scaling exponents for $g_{1}$, $g_{3}$, and $D$ for ideal rings, isolated self-avoiding (S.A.) rings, and rings in the melt state.
$g_{1}$ $g_{3}$ $D$
Ideal rings
Rouse Dynamics for the quenched tree $1/3$ $1$ $-1$
Ring in an array of obstacles $1/5$ $3/5$ $-2$
Ring slithering around the tree $1/4,1/8,1/4$ $1/2$ 00
S.A. rings
Rouse Dynamics for the quenched tree $1/2$ $1$ $-1$
Ring in an array of obstacles $3/8$ $3/4$ $-5/3$
Ring slithering around the tree $1/2,1/4,1/2$ $1/2$ 00
Melt rings
Rouse Dynamics for the quenched tree $2/5$ $1$ $-1$
Ring in an array of obstacles $6/23$ $15/23$ $-17/9$
Ring slithering around the tree $1/3,1/6,1/3$ $1/2$ 00
FLG $2/7$ $5/7$ $-5/3$
II.2.3 Brownian dynamics of randomly quenched trees
@@
The emerging dynamics of quenched trees, where only tree nodes can diffuse via Brownian motion, are expected to be described by a suitably generalized Rouse model [35] predicting free diffusion of the tree CM at all time scales, $g_{3}(t)\sim t/N$, $g_{1}(t)\sim t^{\frac{2\nu}{2\nu+1}}$ in the intermediate regime and $g_{1}(t)\sim t$ on the shortest time scales. Moreover, the CM diffusion coefficient is expected to scale as $D\sim N^{-1}$.
II.2.4 Rings in an array of transient obstacles
@@
The dynamics in the melt state is subject to local topological constraints due to the mutual impenetrability of double-folded ring sections. However, to the extent that the rings move, these constraints are not fixed, but transient. The recent fractal loopy globule (FLG) model, incorporates the effect of moving obstacles on the dynamics of the target ring [36]. According to this model, topological constraints in the melt state gradually dilute over time. This occurs because, as time progresses, larger loops relax, eventually no longer acting as obstacles. This process is analogous to the tube dilation phenomenon observed in polydisperse linear melts [37, 38, 39]. The FLG model predicts the scaling of the mean square displacement as $g_{1}(t)\sim t^{\frac{2\nu}{2+\nu}}$, and diffusion coefficient scaling as $D\sim N^{-2+\nu}$.
III Model and Method
III.1 Discrete Double-Folded Ring Model

### v2

@@ Abstract
The statistics of randomly branching double-folded ring polymers are relevant to the secondary structure of RNA, the large-scale branching of plectonemic DNA (and thus bacterial chromosomes), the conformations of single-ring polymers migrating through an array of obstacles, as well as to the conformational statistics of eukaryotic chromosomes and melts of crumpled, non-concatenated ring polymers. Double-folded rings fall into different dynamical universality classes depending on whether the random tree-like graphs underlying the double-folding are quenched or annealed, and whether the trees can undergo unhindered Brownian motion in their spatial embedding. Locally, one can distinguish (i) repton-like mass transport around a fixed tree, (ii) the spontaneous creation and deletion of side branches, and (iii) displacements of tree nodes, where complementary ring segments diffuse together in space. Here we employ dynamic Monte Carlo simulations of a suitable elastic lattice polymer model of double-folded, randomly branching ring polymers to explore the mesoscopic dynamics that emerge from different combinations of the above local modes in three different systems: ideal non-interacting rings, self-avoiding rings, and rings in the melt state. We observe the expected scaling regimes for ring slithering, the dynamics of double-folded rings in an array of obstacles, and Rouse-like tree dynamics as limiting cases. Notably, the monomer mean-square displacements of $g_{1}\sim t^{0.4}$ observed for crumpled rings with $\nu=1/3$ are similar to the subdiffusive regime observed in bacterial chromosomes. In our analysis, we focus on the question to which extent contributions of different local dynamical modes to the emergent dynamics are simply additive. In particular, we reveal a non-trivial acceleration of the dynamics of interacting rings, when all three types of local motion are present. In the melt case, the asymptotic ring center-of-mass diffusion is dominated by the contribution from coupling of the ring-in-an-array-of-obstacles dynamics with the tree dynamics. This contribution scales inversely with the ring weight and is compatible with a scenario in which constraint release restores a Rouse-like dynamics.
[…]
I.  INTRODUCTION
@@
Large ring polymers exhibit surprisingly rich behavior. When diffusing through an array of obstacles, they naturally adopt double-folded conformations, forming randomly branching polymer structures or trees 1, 2, 3, 4, 5, 6 (Fig. 1). Interestingly, double-folded tree-like ring conformations have broad significance in both physical 7 and biological systems. Biological examples include plectonemes in supercoiled DNA molecules 8, 9, 10 and, consequently, their role in bacterial chromosome structuring 11, 12, 9, 13, large RNA molecules 14, 15, 16, and the arrangement of DNA in eukaryotic cells during interphase, 17, 18, 19, 20 which bears striking similarities to the behavior of melts of non-concatenated ring polymers. 21, 22, 23, 5, 24 Areas of current research 25 include the full characterization of threading phenomena, 26, 24 where inter-ring penetrations impose long-lived topological constraints 27 that dramatically slow down relaxation 28, 29, 30, as well as the alternative scenario of activated dynamics induced by inter-ring caging 31, 32. Yet, the scaling of the center-of-mass diffusion coefficient $D$ with ring length $N$ in melts remains unresolved. None of the current theoretical predictions for the scaling exponent are consistent with molecular dynamics simulations, which observe slower diffusion 23.
@@
Figure 1: (a) An example of a (tightly wrapped) double-folded ring polymer and (b) the corresponding tree on a trigonal lattice. Small loops represent reptons (bonds of zero length, see Section 3), for which adjacent monomers along the ring occupy identical lattice sites.
@@
Here we explore the dynamics of tightly double-folded using a model that does not allow for threading. Double-folded rings fall into different dynamical universality classes depending on whether the random tree-like graphs underlying the double-folding are quenched or annealed 33, and whether the trees can undergo unhindered Brownian motion in their spatial embedding. Specifically, one can identify three distinct types of local dynamics:
@@
    Reptons transporting ring mass along a fixed tree 34, 10 (Fig. 2(a)).
@@
    Creation or retraction of side-branches where either (i) one of the two double-folded ring segments double-folds onto itself and branches away from the current tree or where vpnversly (ii) two antiparallel ring segments wrapping a side branch are retracted into the rest of the tree in the inverse process (Fig. 2(b)).
@@
    Brownian motion of tree nodes maintaining the tree connectivity and the secondary structure of the double-folded rings 35 (Fig. 2(c)-(d)).
The effect of repeated local modifications (or “moves” in the Monte Carlo (MC) language we will adopt throughout the article) is not simply additive to the resulting dynamics of the ring and tree conformations. Instead, an emergent dynamics typically arises because the ring “monomers” and tree “nodes” are not independent degrees of freedom but are coupled through (i) the ring connectivity, (ii) the constraint of double-folding restricting ring conformations to those with underlying tree structures, (iii) additional Hamiltonian terms governing interactions between monomers or nodes, branching activity, or chain stiffness, and (iv) local topological constraints arising from the mutual impenetrability of ring and tree backbones. The emerging internal dynamics typically displays intermediate scaling regimes, whose understanding is important to comprehend the long-time diffusion constant.
This coupling is not a particularity of rings. In linear chains, the connectivity between successive monomers gives rise to the characteristic Rouse intrachain dynamics, while topological constraints are at the origin of the reptation dynamics in the tube model 36. Dynamical analogues can be formulated for rings. In particular, a well-understood limit is the “ring-in-an-array-of-obstacles” (RiAO) dynamics 2, 6, 37, which arises from the coupling of local mass transport along the tree and the creation and deletion of side branches. However, understanding the general case remains a challenge, more particularly in the melt case, where the “obstacles” represent co-localized ring segments that are not fixed in space but instead have a dynamics of their own 37.
The purpose of the present paper is to pave the way for solving this challenge by introducing:
@@
  • a procedure to analyze the independency of contributions from different dynamic modes to the experimentally observable monomer and the center-of-mass mean-square displacements.
The manuscript is structured as follows: In Section 2 we introduce the relevant observables and associated exponents, along with a brief summary of the theoretical background. In Section 3, we introduce our model and the corresponding dynamic MC scheme to investigate its dynamical properties. In particular, this scheme extends our previous elastic lattice model for simulating tightly double-folded ring polymers, where only single monomers could move 38, by incorporating an additional tree node move akin to diffusive dynamics. We then explain how to systematically and in a fully controlled manner combine the different types of monomer and tree node moves. Finally, we introduce the systems studied: ideal non-interacting double-folded rings, isolated self-avoiding double-folded rings, and melts of double-folded rings. In the Results Section 4, we briefly examine the well-understood static ring properties of these systems, which also serve to validate that the ensembles generated by different combinations of local dynamics remain identical as long as the system conditions are unchanged. We then focus on comparing the diffusion properties of both monomers and the center of mass across different dynamical schemes with available theoretical results. Notably, in melts, we find that the mixture of tree node and ring monomer moves leads to a non-trivial acceleration of the emergent dynamics. The subsequent Discussion Section 5 explores the functional form and amplitude of the dynamics arising from the coupling in a context that deals more generally with the analysis of data from mixed dynamical schemes and the question of how to identify and distinguish non-trivial coupling effects from situations where the contributions of different local modes to the emergent dynamics are independent. Finally, we provide a brief conclusion in Section.6.
II.  THEORETICAL BACKGROUND
2.1 Static Properties
The behavior of a double-folded ring polymer 1, 2, 3, 23, 6, 5, 39 can be analyzed at three structural levels 35 (Fig. 1). The primary structure is defined as the connectivity of the monomers along the ring. The secondary structure corresponds to the double folding. It can be represented by mapping the ring onto a graph with the connectivity of the tree. The tertiary structure is defined by the positions of the ring monomers and tree nodes within the embedding (three-dimensional) space. A small set of exponents describes how the expectation values of typical observables scale with the ring weight $N$ (the number of monomers):
@@ 2.1 Static Properties
where $L$ is the mean tree contour distance (the shortest distance on the tree) between tree nodes, $R_{g}$ is the radius of gyration of the ring 38. For ideal, non-interacting systems, the exponents are $\rho=1/2$ and $\nu=1/4$ 40, 41. For interacting systems, the only known exact result is the value $\nu=1/2$ for self-avoiding trees in $d=3$ 42. Otherwise, Flory theory 43, 44, 33, 4, 45 predicts $\rho=2/3$ for self-avoiding trees, and $\nu=1/3$ and $\rho=5/9$ for trees in the melt state.
2.2 Dynamical Properties
@@
On time scales where the rings have moved beyond their own size, their CM exhibit diffusive motion: $g_{1}(t)=g_{3}(t)\sim t$, and $g_{2}(t)=2\langle R^{2}_{g}\rangle$. On shorter time scales, the internal dynamics can display a range of intermediate scaling regimes, whose understanding is important to comprehend the diffusion constant in the asymptotic regime. A summary of the scaling exponents for $g_{1}(t)$, $g_{3}(t)$, and the diffusion constant $D$ in the known limiting cases are discussed below and given in Table 1.
2.2.1 Slithering dynamics of tightly wrapped rings around fixed trees
For local mass transport via reptons along a fixed tree 46, 36, 34, 10, 47, the monomer motion in the embedding space is given by:
@@ 2.2.1 Slithering dynamics of tightly wrapped rings around fixed trees
where $\langle\delta s^{2}(t)\rangle$ denotes the one-dimensional monomer MSD along the tree. Several regimes must then be distinguished. First, for the shortest time scales, $\langle\delta s^{2}(t)\rangle\sim t$, such that:
@@
At longer time scales, the ring connectivity gives rise to one-dimensional Rouse dynamics typical of a linear polymer, $\langle\delta s^{2}(t)\rangle\sim t^{1/2}$, such that:
@@
Beyond the Rouse time, $\tau_{R}\sim N^{2}$ where $g_{1}(\tau_{R})\sim N^{\nu}\sim\sqrt{\langle R_{g}^{2}\rangle}$, the dynamics crosses over to the second reptation regime where all monomers diffuse coherently around the tree. Again $\langle\delta s^{2}(t)\rangle\sim t$, so that:
@@
2.2.2 Rings in an array of fixed obstacles
@@
The combination of repton moves with the local creation or retraction of side branches gives rise to the characteristic motion of rings in an array of fixed obstacles 2, 6 suppressing the sideways motion of tree nodes. Specifically, $g_{1}(t)\sim t^{\frac{2\nu}{2+\rho}}$ and $g_{3}(t)\sim t^{\frac{2\nu+1}{2+\rho}}$ in the intermediate regime 6. Moreover, the CM diffusion coefficient is found to scale as $D\sim N^{-2-\rho+2\nu}$ 6.
@@
Table 1: Summary of the scaling exponents for both static and dynamic properties of ideal rings, isolated self-avoiding (S.A.) rings, and self-avoiding rings in the melt state. Specifically, the table reports the dynamic exponents for $g_{1}(t)$, $g_{3}(t)$, and the diffusion constant $D$, as well as the static exponents $\nu$ and $\rho$.
$\nu$ $\rho$ $g_{1}$ $g_{3}$ $D$
Ideal rings $1/4$ $1/2$
Rouse Dynamics for the quenched tree $1/3$ 11 1-1
Ring in an array of obstacles $1/5$ $3/5$ 2-2
Ring slithering around the tree $1/4,1/8,1/4$ $1/4$ 0
S.A. rings $1/2$ $2/3$
Rouse Dynamics for the quenched tree $1/2$ 11 1-1
Ring in an array of obstacles $3/8$ $3/4$ 5/3-5/3
Ring slithering around the tree $1/2,1/4,1/2$ $1/2$ 0
Melt rings $1/3$ $5/9$
Rouse Dynamics for the quenched tree $2/5$ 11 1-1
Ring in an array of obstacles $6/23$ $15/23$ 17/9-17/9
Ring slithering around the tree $1/3,1/6,1/3$ $1/3$ 0
Fractal Loopy Globule (FLG) $2/7$ $5/7$ 5/3-5/3
2.2.3 Brownian dynamics of randomly quenched trees
@@
The emerging dynamics of quenched trees, where only tree nodes can diffuse via Brownian motion, are expected to be described by a suitably generalized Rouse model 48 predicting free diffusion of the tree CM at all time scales, $g_{3}(t)\sim t/N$, $g_{1}(t)\sim t^{\frac{2\nu}{2\nu+1}}$ in the intermediate regime and $g_{1}(t)\sim t$ on the shortest time scales. Moreover, the CM diffusion coefficient is expected to scale as $D\sim N^{-1}$.
2.2.4 Rings in an array of transient obstacles
@@
The dynamics in the melt state is subject to local topological constraints due to the mutual impenetrability of double-folded ring sections. However, to the extent that the rings move, these constraints are not fixed, but transient. The fractal loopy globule (FLG) model incorporates the effect of moving obstacles on the dynamics of the target ring 37. According to this model, topological constraints in the melt state gradually dilute over time. This occurs because, as time progresses, larger loops relax, eventually no longer acting as obstacles. This process is analogous to the tube dilation phenomenon observed in polydisperse linear melts 49, 50, 51. The FLG model predicts the scaling of the mean square displacement as $g_{1}(t)\sim t^{\frac{2\nu}{2+\nu}}$, and the diffusion coefficient scaling as $D\sim N^{-2+\nu}$.
III.  MODEL AND METHOD
3.1 Discrete Double-Folded Ring Model